# Learning to Steer Markovian Agents under Model Uncertainty

### v2

@@ Abstract
Designing incentives for an adapting population is a ubiquitous problem in a wide array of economic applications and beyond. In this work, we study how to design additional rewards to steer multi-agent systems towards desired policies without prior knowledge of the agents’ underlying learning dynamics. Motivated by the limitation of existing works, we consider a new and general category of learning dynamics called Markovian agents. We introduce a model-based non-episodic Reinforcement Learning (RL) formulation for our steering problem. Importantly, we focus on learning a history-dependent steering strategy to handle the inherent model uncertainty about the agents’ learning dynamics. We introduce a novel objective function to encode the desiderata of achieving a good steering outcome with reasonable cost. Theoretically, we identify conditions for the existence of steering strategies to guide agents to the desired policies. Complementing our theoretical contributions, we provide empirical algorithms to approximately solve our objective, which effectively tackles the challenge in learning history-dependent strategies. We demonstrate the efficacy of our algorithms through empirical evaluations [^1] The code of all the experiments in this paper can be found in https://github.com/jiaweihhuang/Steering_Markovian_Agents..
@@ 1 Introduction
To our knowledge, Zhang et al., (2023) is the first work studying a similar steering problem as ours. They assume the agents are no-regret learners and may act in arbitrarily adversarial ways. In some natural settings, such assumption may not be practical, because no-regret criterion typically requires careful processing of the entire history of learning. In settings with limited cognitive resources and bounded rationality, it is natural to favor models where the agents only process a subset of the available information (Camerer,, 2011). In particular, humans have been widely shown to rely overproportionally on recent experiences in decision making, known as ‘recency bias’ (Costabile and Klein,, 2005; Page and Page,, 2010; Durand et al.,, 2021). Besides, there is evidence that behavioral dynamics that only rely on the most recent experience are able to fit behavioral data well in certain situations (Mäs and Nax,, 2016). Motivated by these insights, we therefore study steering a different category of learning dynamics called Markovian agents, where the agents’ policy updates only depend on their current policy and the (modified) reward function. Our model complements the prior work on no-regret agents, and serves as the first abstraction of behavior based on limited cognitive abilities with recency bias in steering setting. Theoretically, Markovian agents subsumes a broader class of popular policy-based methods as concrete examples (Giannou et al.,, 2022; Ding et al.,, 2022), which are not covered by no-regret assumptions. We also note that a concurrent work (Canyakmaz et al.,, 2024) considers a similar setting as ours, and we defer to Sec. 1.1 for further discussion.
@@ 1.1 Closely Related Works
Perhaps the closest to ours is a concurrent work by Canyakmaz et al., (2024), which contributes empirical investigation on the use of control methods to direct game dynamics towards desired outcomes, in particular allowing for model uncertainty. Although not formally specified in Canyakmaz et al., (2024), they consider a similar finite-horizon non-episodic setup as ours. However, they only consider history-independent steering strategy, which can result in sub-optimal performance in this finite-horizon setup, because one-step observations may not provide sufficient information to tackle model uncertainty. As our main contribution compared with their work, we point out that, one should, in principle, employ history-dependent steering strategies, since history can serve as sufficient information set for decision making under uncertainty. This leads to significant differences in the design principles of our algorithms compared with Canyakmaz et al., (2024). Concretely, we propose a learning objective for history-dependent strategies in Obj. (1), and two algorithms for low uncertainty (small $\mathcal{F}$) and high uncertainty (large $\mathcal{F}$) settings, respectively. In the former case, we contribute a belief-state based algorithm that can exactly solve Obj. (1), offering a stronger solution than Canyakmaz et al., (2024) due to the theoretical guarantee in Prop. 3.3. For the latter, although both our FETE and SIAR-MPC (Canyakmaz et al.,, 2024) share a two-phase (exploration + exploitation) structure, ours represents a more general framework with a more advanced exploration strategy (see more explanation in Sec. 5.2). Besides, we develop additional novel theory regarding the existence of strategies with low steering gap.
@@ 3 The Problem Formulation of the Steering Markovian Agents
Here $\psi:=\{\psi_{t}\}_{t\in[T]}$ denotes the mediator’s “steering strategy” to generate $u_{t}$. We consider history-dependent strategies to handle the model uncertainty, which we will explain later. Besides, ${\bm{u}}_{t}:=\{u_{t,h}^{n}\}_{h\in[H],n\in[N]}$, where $u^{n}_{t,h}:\mathcal{S}\times\mathcal{A}\rightarrow[0,U_{\max}]$ is the steering reward for agent $n$ at game horizon $h$ and steering step $t$. $U_{\max}<+\infty$ denotes the upper bound for the steering reward. For practical concerns, we follow the standard constraints that the steering rewards are non-negative.
@@
$f^{*}\in\mathcal{F}$.
A Finite-Hoziron Non-Episodic Setup and Motivation As motivated previously, we formulate steering as a finite-horizon non-episodic RL problem. To our knowledge, in contrast to our finite-horizon setting, most of the non-episodic RL settings consider the infinite-horizon setup with stationary or non-stationary transitions, and therefore, they are also not suitable here. We provide more discussion in Sec. 1.1.
@@
The mediator can only interact with the real agents for one episode $\{{\bm{\pi}}_{1},{\bm{u}}_{1},...,{\bm{\pi}}_{T},{\bm{u}}_{T},{\bm{\pi}}_{T+1}\}$, where ${\bm{\pi}}_{t+1}\sim f^{*}(\cdot|{\bm{\pi}}_{t},{\bm{u}}_{t})~{}\forall t\in[T]$. Nonetheless, the mediator can get access to the simulators for all models in $\mathcal{F}$, and it can sample arbitrary trajectories and do episodic learning with those simulators to decide the best steering actions ${\bm{u}}_{1},{\bm{u}}_{2},...,{\bm{u}}_{T}$ to deploy.
@@
We want to clarify that in our steering framework, we will first solve Obj. (1), and then deploy $\psi^{*}$ to steer real agents. The learning and optimization of $\psi^{*}$ in Obj. (1) only utilizes simulators of $\mathcal{F}$. Besides, after deploying $\psi^{*}$ to real agents, we will not update $\psi^{*}$ with the data generated during the interaction with real agents. This is seemingly different from common online learning algorithms which conduct the learning and interaction repeatedly(Dann and Brunskill,, 2015). But we want to highlight that, given the fact that $\psi^{*}$ is history-dependent, it is already encoded in $\psi^{*}$ how to make decisions (or say, learning) in the face of uncertainty after gathering data from real agents. In other words, one can interpret that, in Obj. (1), we are trying to optimize an “online algorithm” $\psi^{*}$ which can “smartly” decide the next steering reward to deploy given the past interaction history. As we will justify in the following, our Obj. (1) can indeed successfully handle the model uncertainty.
Justification for Objective (1) We use $C_{\psi,T}(f):=\mathbb{E}_{\psi,f}[\sum_{t=1}^{T}\eta^{\text{cost}}({\bm{\pi}}_{t},{\bm{u}}_{t})]$ and $\Delta_{\psi,T}(f):=\mathbb{E}_{\psi,f}[\max_{\bm{\pi}}\eta^{\text{goal}}({\bm{\pi}})-\eta^{\text{goal}}({\bm{\pi}}_{T+1})]$ as short notes of the steering cost and the steering gap (of the terminal policy ${\bm{\pi}}_{T+1}$), respectively. Besides, we denote $\Psi^{\varepsilon}:=\{\psi\in\Psi|\max_{f\in\mathcal{F}}\Delta_{\psi,T}(\mathcal{F})\leq\varepsilon\}$ [^3] In fact, besides $\varepsilon$, $\Psi^{\varepsilon}$ also depends on other parameters like $T$, $U_{\max}$, $\mathcal{F}$ and the initial policy ${\bm{\pi}}_{1}$. For simplicity, we only highlight those dependence if necessary. to be the collection of all steering strategies with $\varepsilon$-steering gap. Based on these notations, we introduce two desiderata, and show how an optimal solution $\psi^{*}$ of Obj. (1) can achieve them.
@@ 4 Existence of Steering Strategy with $\varepsilon$-Steering Gap
In this section, we identify sufficient conditions such that $\Psi^{\varepsilon}$ is non-empty. In Sec. 4.1, we start with the special case when $f^{*}$ is known, i.e. =$\mathcal{F}=\{f^{*}\}$. The results will serve as basis when we study the general unknown model setting in Sec. 4.2.
@@ 4.1 Existence when $f^{*}$ is Known: Natural Policy Gradient as an Example
For NPG, note that the policy is always bounded away from 0. We will use $\Pi^{+}:=\{{\bm{\pi}}|\forall n,h,a_{h},s_{h}:\pi^{n}_{h}(a_{h}|s_{h})>0\}$ to denote such feasible policy set. We state our main result below.
@@
Suppose $\eta^{\text{goal}}$ is Lipschitz in ${\bm{\pi}}$, given any initial ${\bm{\pi}}_{1}\in\Pi^{+}$, for any $\varepsilon>0$, if the agents follow Def. 4.1 under Assump. B, if $T$ and $U_{\max}$ are large enough, we have $\Psi^{\varepsilon}\neq\emptyset$.
@@ 4.2 Existence when $f^{*}$ is Unknown: the Identifiable Model Class
Intuitively, $\mathcal{F}$ is $(\delta,T_{\mathcal{F}}^{\delta})$-identifiable, if $\exists\psi$, s.t. after $T_{\mathcal{F}}^{\delta}$ steering steps, the hidden model $f$ can be identified by the Maximal Likelihood Estimation (MLE) with high probability. Next, we provide an example of $(\delta,T_{\mathcal{F}}^{\delta})$-identifiable function class with $T_{\mathcal{F}}^{\delta}$ upper bounded for any $\delta\in(0,1)$.
@@
[A Sufficient Condition for Existence] Given any $\varepsilon>0$, $\Psi^{\varepsilon}_{T}(\mathcal{F};{\bm{\pi}}_{1})\neq\emptyset$ [^4] Here we highlight the dependence on initial policy, model, and time for clarity (see Footnote 3), if $\exists~{}\widetilde{T}<T$, s.t., (1) $\mathcal{F}$ is $(\frac{\varepsilon}{2\eta_{\max}},\widetilde{T})$-identifiable, (2) $\Psi^{\varepsilon/2}_{T-\widetilde{T}}(\mathcal{F};{\bm{\pi}}_{\widetilde{T}})\neq\emptyset$ for any possible ${\bm{\pi}}_{\widetilde{T}}$ generated at step $\widetilde{T}$ during the steering.
@@ 5.2 Large Model Class: A First-Explore-Then-Exploit Framework
Justification for FETE We cannot guarantee that Desiderata 12 are achievable, because we do not exactly solve Obj. 1. However, if $\mathcal{F}$ is $({\delta}/|\mathcal{F}|,T_{\mathcal{F}}^{\delta/|\mathcal{F}|})$-identifiable (Def. 4.3) and we choose $\widetilde{T}\geq T_{\mathcal{F}}^{\delta/|\mathcal{F}|}$, we can verify $\Pr(f_{\text{MLE}}=f^{*})\geq 1-\delta$ in Proc. 2. Therefore, we can still expect the exploitation policy $\psi^{\text{Exploit}}$ steer the agents to approximately maximize $\eta^{\text{goal}}({\bm{\pi}}_{T+1})$ with reasonable steering cost for the rest $T-\widetilde{T}$ steps.
@@
3 Learn an exploration strategy
$\psi^{\text{Explore}}\leftarrow\arg\max_{\psi}\frac{1}{|\mathcal{F}|}\sum_{f\in\mathcal{F}}\mathbb{E}_{{\bm{\pi}}^{\prime}_{1},{\bm{u}}_{1}^{\prime},...,{\bm{\pi}}^{\prime}_{\widetilde{T}+1}\sim\psi,f}[\mathbb{I}[f=\operatorname*{arg\,max}_{f^{\prime}\in\mathcal{F}}\sum_{t=1}^{\widetilde{T}}\log f^{\prime}({\bm{\pi}}_{t+1}^{\prime}|{\bm{\pi}}^{\prime}_{t},{\bm{u}}^{\prime}_{t})]].$
@@ 3 Learn an exploration strategy
Comparison with Canyakmaz et al., (2024) Although both SIAR-MPC in Canyakmaz et al., (2024) and our FETE (Procedure 2) adopt a first-explore-then-exploit structure, FETE is more general and more effective. Both algorithms have three main components: exploration strategy, model estimation strategy and exploitation strategy, and we inspect our advantages from these three aspects. (i) Exploration strategy: SIAR-MPC uses noise-based random exploration, whereas we adopt a more strategic approach, which uses the identification success rate as a signal to learn the exploration policy. Empirical results in Sec. 6.2 demonstrate the higher efficiency of our methods. (ii) Model estimation strategy: SIAR-MPC estimates the hidden model by solving a regression problem with constraints (Eq. (8) in Canyakmaz et al., (2024)), while we solve a MLE objective. In fact, our MLE objective is more general and can recover the regression problem in SIAR-MPC, if we consider a model class $\mathcal{F}$ that includes Gaussian noise perturbed dynamics with the side-information constraints introduced in Canyakmaz et al., (2024). (iii) Exploitation strategy: As a general framework, our FETE does not restrict how to compute the exploitation strategy $\psi^{Exploit}$. As we suggest in paper, any RL or control method can be used, including the MPC approach in Canyakmaz et al., (2024).

### v3

@@ Abstract
Designing incentives for an adapting population is a ubiquitous problem in a wide array of economic applications and beyond. In this work, we study how to design additional rewards to steer multi-agent systems towards desired policies without prior knowledge of the agents’ underlying learning dynamics. Motivated by the limitation of existing works, we consider a new and general category of learning dynamics called Markovian agents. We introduce a model-based non-episodic Reinforcement Learning (RL) formulation for our steering problem. Importantly, we focus on learning a history-dependent steering strategy to handle the inherent model uncertainty about the agents’ learning dynamics. We introduce a novel objective function to encode the desiderata of achieving a good steering outcome with reasonable cost. Theoretically, we identify conditions for the existence of steering strategies to guide agents to the desired policies. Complementing our theoretical contributions, we provide empirical algorithms to approximately solve our objective, which effectively tackles the challenge in learning history-dependent strategies. We demonstrate the efficacy of our algorithms through empirical evaluations.
@@ 1 Introduction
To our knowledge, Zhang et al., (2023) is the first work studying a similar steering problem as ours. They assume the agents are no-regret learners and may act in arbitrarily adversarial ways. In some natural settings, such assumption may not be practical, because no-regret criterion typically requires careful processing of the entire history of learning. In settings with limited cognitive resources and bounded rationality, it is natural to favor models where the agents only process a subset of the available information (Camerer,, 2011). In particular, humans have been widely shown to rely overproportionally on recent experiences in decision making, known as ‘recency bias’ (Costabile and Klein,, 2005; Page and Page,, 2010; Durand et al.,, 2021). Besides, there is evidence that behavioral dynamics that only rely on the most recent experience are able to fit behavioral data well in certain situations (Mäs and Nax,, 2016). Motivated by these insights, we therefore study steering a different category of learning dynamics called Markovian agents, where the agents’ policy updates only depend on their current policy and the (modified) reward function. Our model complements the prior work on no-regret agents, and serves as the first abstraction of behavior based on limited cognitive abilities with recency bias in steering setting. Theoretically, Markovian agents subsumes a broader class of popular policy-based methods as concrete examples (Giannou et al.,, 2022; Ding et al.,, 2022), which are not covered by no-regret assumptions. We note that a concurrent work (Canyakmaz et al.,, 2024) considers a similar setting as ours, and we defer to Sec. 1.1 for further discussion.
@@ 1.1 Closely Related Works
Perhaps the closest to ours is a concurrent work by Canyakmaz et al., (2024). They consider a similar finite-horizon non-episodic setup as ours, and experimentally investigate the use of control methods to direct game dynamics towards desired outcomes, in particular allowing for model uncertainty. Compared to their work, we handle the model uncertainty in a more principled way by employing history-dependent steering strategies, since history can serve as sufficient information set for decision making under uncertainty. This leads to several differences in the design principles between our algorithms and Canyakmaz et al., (2024). Concretely, we propose a learning objective for history-dependent strategies in Obj. (1), and two algorithms for low uncertainty (small $\mathcal{F}$) and high uncertainty (large $\mathcal{F}$) settings, respectively. In the former case, we contribute a belief-state based algorithm that can exactly solve Obj. (1), offering a stronger solution due to the theoretical guarantee in Prop. 3.3. For the latter, although both our FETE and SIAR-MPC (Canyakmaz et al.,, 2024) share a two-phase (exploration + exploitation) structure, ours represents a more general framework with a more advanced exploration strategy (see more explanation in Sec. 5.2). Besides, we develop additional novel theory regarding the existence of strategies with low steering gap.
@@ 3 The Problem Formulation of the Steering Markovian Agents
Here $\psi:=\{\psi_{t}\}_{t\in[T]}$ denotes the mediator’s “steering strategy” to generate $u_{t}$. We consider history-dependent strategies to handle the model uncertainty, which we will explain later. Besides, ${\bm{u}}_{t}:=\{u_{t,h}^{n}\}_{h\in[H],n\in[N]}$, where $u^{n}_{t,h}:\mathcal{S}\times\mathcal{A}\rightarrow[0,U_{\max}]$ is the steering reward for agent $n$ at game horizon $h$ and steering step $t$. $U_{\max}<+\infty$ denotes the upper bound for the steering reward. With a bit of abuse of notation, here we treat both ${\bm{r}}$ and ${\bm{u}}_{t}$ as vectors with length $HN|\mathcal{S}||\mathcal{A}|$ and use ${\bm{r}}+{\bm{u}}_{t}$ to denote elementwise addition. For practical concerns, we follow the standard constraints that the steering rewards are non-negative.
@@
The true learning dynamics $f^{*}$ is realizable, i.e. $f^{*}\in\mathcal{F}$.
A Finite-Hoziron Non-Episodic Setup and Motivation As motivated previously, we formulate steering as a finite-horizon non-episodic RL problem. To our knowledge, in contrast to our finite-horizon setting, most of the non-episodic RL settings consider the infinite-horizon setup with stationary or non-stationary transitions, and therefore, they are also not suitable here. We provide more discussion in Appx. B.2.
@@
The mediator can only interact with the real agents for one episode $\{{\bm{\pi}}_{1},{\bm{u}}_{1},...,{\bm{\pi}}_{T},{\bm{u}}_{T},{\bm{\pi}}_{T+1}\}$, where ${\bm{\pi}}_{t+1}\sim f^{*}(\cdot|{\bm{\pi}}_{t},{\bm{r}}+{\bm{u}}_{t})$ for all $t\in[T]$. Nonetheless, the mediator can get access to the simulators for all models in $\mathcal{F}$, and it can sample arbitrary trajectories and do episodic learning with those simulators to decide the best steering actions ${\bm{u}}_{1},{\bm{u}}_{2},...,{\bm{u}}_{T}$ to deploy.
@@
We want to clarify that in our steering framework, we will first solve Obj. (1), and then deploy $\psi^{*}$ to steer real agents. The learning and optimization of $\psi^{*}$ in Obj. (1) only utilizes simulators of $\mathcal{F}$. Besides, after deploying $\psi^{*}$ to real agents, we will not update $\psi^{*}$ with the data generated during the interaction with real agents. This is seemingly different from common online learning algorithms which conduct the learning and interaction repeatedly(Dann and Brunskill,, 2015). But we want to highlight that, given the fact that $\psi^{*}$ is history-dependent, it is already encoded in $\psi^{*}$ how to make decisions (or say, learning) in the face of uncertainty after gathering data from real agents. In other words, one can interpret that, in Obj. (1), the $\psi^{*}$ to be optimized is an “online algorithm” which can “smartly” decide the next steering reward to deploy given the past interaction history. As we will justify in the following, our Obj. (1) can indeed successfully handle the model uncertainty.
Justification for Objective (1) We use $C_{\psi,T}(f):=\mathbb{E}_{\psi,f}[\sum_{t=1}^{T}\eta^{\text{cost}}({\bm{\pi}}_{t},{\bm{u}}_{t})]$ and $\Delta_{\psi,T}(f):=\mathbb{E}_{\psi,f}[\max_{\bm{\pi}}\eta^{\text{goal}}({\bm{\pi}})-\eta^{\text{goal}}({\bm{\pi}}_{T+1})]$ as short notes of the steering cost and the steering gap (of the terminal policy ${\bm{\pi}}_{T+1}$), respectively. Besides, we denote $\Psi^{\varepsilon}:=\{\psi\in\Psi|\max_{f\in\mathcal{F}}\Delta_{\psi,T}(\mathcal{F})\leq\varepsilon\}$ [^2] In fact, besides $\varepsilon$, $\Psi^{\varepsilon}$ also depends on other parameters like $T$, $U_{\max}$, $\mathcal{F}$ and the initial policy ${\bm{\pi}}_{1}$. For simplicity, we omit those dependence unless necessary. to be the collection of all steering strategies with $\varepsilon$-steering gap. Based on these notations, we introduce two desiderata, and show how an optimal solution $\psi^{*}$ of Obj. (1) can achieve them.
@@ 4 Existence of Steering Strategy with $\varepsilon$-Steering Gap
In this section, we identify sufficient conditions such that $\Psi^{\varepsilon}$ is non-empty. In Sec. 4.1, we start with the special case when $f^{*}$ is known, i.e. =$\mathcal{F}=\{f^{*}\}$. The results will serve as basis when we study the general unknown model setting with $|\mathcal{F}|>1$ in Sec. 4.2.
@@ 4.1 Existence when $f^{*}$ is Known: Natural Policy Gradient as an Example
Note that the exact NPG is captured by the special case with $\lambda_{\min}=\lambda_{\max}=1$. For NPG, since the policy is always bounded away from 0. We will use $\Pi^{+}:=\{{\bm{\pi}}|\forall n,h,a_{h},s_{h}:\pi^{n}_{h}(a_{h}|s_{h})>0\}$ to denote such feasible policy set. We state our main result below.
@@
Suppose $\eta^{\text{goal}}$ is Lipschitz in ${\bm{\pi}}$, given any initial ${\bm{\pi}}_{1}\in\Pi^{+}$, for any $\varepsilon>0$, if the agents follow Def. 4.1 under Assump. B, and $T$ and $U_{\max}$ are large enough, we have $\Psi^{\varepsilon}\neq\emptyset$.
@@ 4.2 Existence when $f^{*}$ is Unknown: the Identifiable Model Class
Intuitively, $\mathcal{F}$ is $(\delta,T_{\mathcal{F}}^{\delta})$-identifiable, if $\exists\psi$, s.t. after $T_{\mathcal{F}}^{\delta}$ steering steps, the hidden model $f$ can be identified by the Maximal Likelihood Estimation (MLE) with high probability. Next, we provide a general set of $(\delta,T_{\mathcal{F}}^{\delta})$-identifiable function classes with $T_{\mathcal{F}}^{\delta}$ upper bounded for any $\delta\in(0,1)$.
@@
[A Sufficient Condition for Existence] Given any $\varepsilon>0$, $\Psi^{\varepsilon}_{T}(\mathcal{F};{\bm{\pi}}_{1})\neq\emptyset$ [^3] Here we highlight the dependence on initial policy, model, and time for clarity (see Footnote 2), if $\exists~{}\widetilde{T}<T$, s.t., (1) $\mathcal{F}$ is $(\frac{\varepsilon}{2\eta_{\max}},\widetilde{T})$-identifiable, (2) $\Psi^{\varepsilon/2}_{T-\widetilde{T}}(\mathcal{F};{\bm{\pi}}_{\widetilde{T}})\neq\emptyset$ for any possible ${\bm{\pi}}_{\widetilde{T}}$ generated at step $\widetilde{T}$ during the steering.
@@ 5.2 Large Model Class: A First-Explore-Then-Exploit Framework
Justification for FETE We cannot guarantee that Desiderata 12 are achievable, because we do not exactly solve Obj. 1. However, if $\mathcal{F}$ is $({\delta}/|\mathcal{F}|,T_{\mathcal{F}}^{\delta/|\mathcal{F}|})$-identifiable (Def. 4.3) and we choose $\widetilde{T}\geq T_{\mathcal{F}}^{\delta/|\mathcal{F}|}$, we can verify $\Pr(f_{\text{MLE}}=f^{*})\geq 1-\delta$ in FETE. Therefore, we can still expect the exploitation policy $\psi^{\text{Exploit}}$ steer the agents to approximately maximize $\eta^{\text{goal}}({\bm{\pi}}_{T+1})$ with reasonable steering cost for the rest $T-\widetilde{T}$ steps.
@@
3 Learn an exploration strategy $\psi^{\text{Explore}}\leftarrow\arg\max_{\psi}\frac{1}{|\mathcal{F}|}\sum_{f\in\mathcal{F}}\mathbb{E}_{{\bm{\pi}}^{\prime}_{1},{\bm{u}}_{1}^{\prime},...,{\bm{\pi}}^{\prime}_{\widetilde{T}+1}\sim\psi,f}[\mathbb{I}[f=\operatorname*{arg\,max}_{f^{\prime}\in\mathcal{F}}\sum_{t=1}^{\widetilde{T}}\log f^{\prime}({\bm{\pi}}_{t+1}^{\prime}|{\bm{\pi}}^{\prime}_{t},{\bm{u}}^{\prime}_{t})]].$
@@
Comparison with Canyakmaz et al., (2024) Both SIAR-MPC in Canyakmaz et al., (2024) and our FETE (Procedure 2) adopt a first-explore-then-exploit structure. We examine the differences between two algorihms from two aspects: exploration strategy and model estimation strategy. Regarding exploration strategy, SIAR-MPC uses noise-based random exploration, whereas we adopt a more strategic approach, which uses the identification success rate as a signal to learn the exploration policy. Empirical results in Sec. 6.2 demonstrate the higher efficiency of our methods. In terms of model estimation strategy, SIAR-MPC estimates the hidden model by solving a constrained regression problem (Eq. (8) in Canyakmaz et al., (2024)), while we solve a MLE objective. In principle, our MLE objective is more general and can recover the regression problem in SIAR-MPC if choosing $\mathcal{F}$ to be the function class including Gaussian noise perturbed dynamics with side-information constraints introduced in Canyakmaz et al., (2024).